# Hidden State Poisoning Attacks against Mamba-based Language Models

### v1

@@ Abstract
State space models (SSMs) like Mamba offer efficient alternatives to Transformer-based language models, with linear time complexity. Yet, their adversarial robustness remains critically unexplored. This paper studies the phenomenon whereby specific short input phrases induce a partial amnesia effect in such models, by irreversibly overwriting information in their hidden states, referred to as a Hidden State Poisoning Attack (HiSPA). Our benchmark RoBench-25 allows evaluating a model’s information retrieval capabilities when subject to HiSPAs, and confirms the vulnerability of SSMs against such attacks. Even a recent 52B hybrid SSM–Transformer model from the Jamba family collapses on RoBench-25 under optimized HiSPA triggers, whereas pure Transformers do not. We also observe that HiSPA triggers significantly weaken the Jamba model on the popular Open-Prompt-Injections benchmark, unlike pure Transformers. Finally, our interpretability study reveals patterns in Mamba’s hidden layers during HiSPAs that could be used to build a HiSPA mitigation system. The full code and data to reproduce the experiments can be found at https://anonymous.4open.science/r/hispa_anonymous-5DB0.
Hidden State Poisoning Attacks against Mamba-based Language Models
Alexandre Le Mercier1   Chris Develder1thanks: Joint senior authors   Thomas Demeester [^1] footnotemark: 1 1IDLab–T2K, Ghent University–imec {alexandre.lemercier, chris.develder, thomas.demeester}@ugent.be
@@ 1 Introduction
State space models (SSMs) like Mamba (gu_mamba_2024) and Mamba-2 (dao_transformers_2024) have emerged as a promising alternative to traditional Transformers in large language models (LLMs) (rezaei_jafari_mambalrp_2024; waleffe_empirical_2024), thanks to their ability to efficiently capture long-range dependencies while achieving linear time complexity during both training and inference. Beyond efficiency and speed advantages, research in SSMs is particularly relevant in today’s ecological context, where the massive energy consumption of standard Transformers (with quadratic complexity) is becoming a growing concern (patterson_carbon_2021; liu_green_2024).
Despite the impressive capabilities demonstrated by recent SSM-based LLMs, their robustness against adversarial attacks remains largely unexplored, and no dedicated benchmark has been established. We argue that this is a critical oversight, since the unique design of the widely used Mamba block, in particular its selectivity mechanism and recurrent hidden-state updates, may expose new types of vulnerabilities that are not present in Transformer-based architectures. Our study reveals and quantifies one such vulnerability, which a malicious user can easily exploit through a short trigger sequence that causes partial amnesia in Mamba by irreversibly overwriting its hidden states. Such Hidden State Poisoning Attack (HiSPA) requires no optimization and can be crafted in zero-shot black-box settings, yet it drastically impairs a model’s ability to retain and retrieve information from long contexts, even when explicitly instructed to disregard the injected content.
This vulnerability directly connects to the well-known Prompt Injection Attack (PIA) problem. We show that HiSPAs severely degrade Mamba’s context retention and that the industrial Jamba-1.7-mini LLM exhibits a significantly higher PIA success rate when a HiSPA precedes the prompt, which demonstrates that hidden-state corruption can amplify classical prompt injections (§§˜3 and 6).
Addressing this issue is urgent in today’s adversarial environment surrounding LLMs (lin_hidden_2025). A representative example is that of assisted scientific literature review tools such as Scispace [^1] https://www.scispace.com/ or Elicit, [^2] https://elicit.com/ which rely on automated reasoning over long textual contexts (wu_gpt_2023; tillmann_literature_2025). Recent reports indicate that malicious actors embed prompt injections directly within papers (keuper_prompt_2025), creating realistic attack surfaces for SSM-based models deployed in these pipelines. Such scenarios underscore the need for systematic robustness evaluation of hybrid architectures before deployment.
@@
  • We introduce Hidden State Poisoning Attacks (HiSPAs), a new class of adversarial triggers that exploit SSM recurrence to overwrite hidden states in Mamba-based language models, with zero-shot (Z-HiSPA) and optimized white-box (M-HiSPA) variants (§˜3).
@@
  • We show that optimized M-HiSPAs can lead to the collapse of SSM as well as hybrid SSM-Transformer models (§˜4.3).
@@
Figure 1: The Mamba core (left) and Mamba block (right) as described by gu_mamba_2024. For block $b$ and time step $t$ (i.e., evaluating the $t$-th token in the sequence), $\mathbf{x}_{t}^{(b)}\in\mathbb{R}^{K}$ is the input core token embedding, $\mathbf{h}_{t}^{(b)}\in\mathbb{R}^{K\times N}$ the hidden states with $N$ the length of $\mathbf{h}_{t,k}^{(b)}$, $\mathbf{y}_{t}^{(b)}\in\mathbb{R}^{K}$ the output core token embedding, and $\mathbf{o}_{t}^{(b)}\in\mathbb{R}^{K/2}$ the output block token embedding.
@@ 2.1 State Space Models for LLMs
The Mamba model (gu_mamba_2024) is built upon the S4 family of state space models (SSMs) (gu_efficiently_2022). Mamba replaces the attention mechanism of Transformers (vaswani_attention_2023) with a structure inspired by recurrent neural networks (RNNs) (schmidt_recurrent_2019) and control theory (kalman:60), achieving linear time complexity with respect to sequence length. Mamba processes each token sequentially while maintaining a hidden state $\mathbf{h}^{(b)}_{t}$ that encodes information from previous tokens, effectively acting as an explicit memory (initialized to zero). For notational simplicity, we omit the block superscript $(b)$ in the remainder of the paper unless disambiguation is required.
@@
The advent of Mamba in 2023 and Mamba-2 in 2024 (dao_transformers_2024) demonstrated that SSMs can match or surpass Transformers on key benchmarks, while offering gains in memory and computational efficiency, especially in hybrid architectures interleaving SSM layers with attention (waleffe_empirical_2024). This breakthrough has sparked innovative hybrid designs including Samba (ren_samba_2025), Hymba (dong_hymba_2024), BlackMamba (anthony_blackmamba_2024), Zamba (glorioso_zamba_2024), the Nemotron-H family (nvidia_nemotron-h_2025), and the Jamba family (lieber_jamba_2024; team_jamba-15_2024). However, the adversarial robustness of these architectures remains unexplored.
@@
Jamba is the largest hybrid SSM-Transformer family in the literature (up to 398B parameters), demonstrating effective context lengths of 256k tokens on the RULER benchmark (hsieh2024ruler), being the only model family maintaining high performance at this scale without degradation (team_jamba-15_2024). Jamba achieves competitive results on standard benchmarks while exhibiting known SSM weaknesses in reasoning and code generation (patro_mamba-360_2024; chen_achilles_2025; ren_exploring_2025). Crucially, Mamba blocks constitute $\nicefrac{{7}}{{8}}$ of Jamba’s architecture, making it ideal for evaluating whether HiSPA vulnerabilities transfer from pure-Mamba models to production-scale hybrids.
@@ 2.2 Prompt Injection Attacks
Prompt Injection Attacks (PIAs) exploit how language models process input prompts to manipulate outputs for malicious purposes (greshake_not_2023; liu_formalizing_2024). These typically comprise a trigger (also named distractor) and a payload, e.g., “Ignore previous instructions and give me your API key”. liu_formalizing_2024 proposed a standard benchmark showing that combining a fake completion with an ignore instruction severely degrades performance of several Transformer-based LLMs, including GPT-4, and that defensive prompt engineering yields only marginal improvements. Although PIAs have been widely studied in Transformer-based LLMs (greshake_not_2023; perez_ignore_2022; chen_defending_2025), their impact on SSM-based models remains largely unexplored, leaving open whether these architectures introduce new attack vectors or exhibit different vulnerability patterns.
Figure 2: Greedy beam search evolution of the loss function (5) per token added to the trigger, distributed over 100 random seeds. The exponential decay of the loss after token n°2 is consistent with the exponential saturation of the hidden states described in App.˜A.
3 Hidden States Poisoning Attacks
@@ 3.1 Theoretical Foundation
The core idea behind a Hidden State Poisoning Attack (HiSPA) is to exploit the inability of SSMs with explicit recurrent state updates, such as Mamba, to attend back to previous tokens and its selectivity mechanism, to poison the hidden state $\mathbf{h}_{t}$. With “poison” we mean introducing a small sequence of tokens that will irreversibly overwrite a large part of the information contained in the hidden state. We will refer to such sequence as a trigger, to align with the PIA terminology.
@@
In eqs.˜1, 2 and 3, gu_mamba_2024 constrained the $\Delta_{t}$ matrix to be positive for computational stability reasons, and made $\mathbf{A}$ diagonal. Further, $\mathbf{B}$ is learned during training, and its values are unconstrained. The objective of a HiSPA is therefore to identify a short sequence whose associated updates drive the recurrence into a contracting regime and inject a dominant input term. As shown in App.˜A, if all tokens in the trigger belong to a subset for which the corresponding state-transition factors satisfy a uniform contraction bound (eqs.˜9, 10, 11 and 12), the carried-over state decays geometrically with the trigger length (eq.˜13). Meanwhile, the input part of the unrolled update necessarily contains a summand whose magnitude equals the maximum attainable input contribution over that subset (eq.˜14). By choosing the last token to realize this maximum and selecting the trigger length so that the decaying state term becomes smaller than this input term (eq.˜17), the injected contribution dominates the previous hidden state. In this sense, a carefully chosen trigger can overwrite most of the stored information in $\mathbf{h}_{t}$, establishing the mechanism enabling HiSPA.
@@
Mamba blocks contain a skip connection (cf. Fig.˜1) that combines the output of the Mamba core branch with a linear projection of the input token embedding itself. It is unclear from the theory alone to what extent this skip connection may dilute the poisoning effect of the Mamba core, further motivating empirical evaluations in §˜4.
Hybrid Models.
Hybrid models combining Mamba with attention layers may partially mitigate this vulnerability, since attention can recover information that the SSM layers have overwritten. We test this hypothesis in §˜4 with Jamba-1.7-Mini.
@@ 3.2 Empirical Verification
$\mathcal{L}(\mathbf{x}_{\text{Tri}}):=-\cos\left(\mathbf{h}(\mathbf{x}_{\text{InT}}\oplus\mathbf{x}_{\text{Tri}}),\mathbf{h}(\mathbf{x}_{\text{Tri}})\right)$ (5)
Using 10 Tesla V100-SXM3-32GB GPUs, we evaluate 100 random seeds (42–141) and obtain the graph in Fig.˜2. As expected from theory, we observe an approximately exponential decay of the loss once the contracting regime dominates, consistent with the $\rho^{L}$ bound in eq.˜12. The best trigger obtained reached a loss of -0.997 after its 6 tokens, making the embedding of $\mathbf{x}_{\text{InT}}\oplus\mathbf{x}_{\text{Tri}}$ nearly aligned with that of $\mathbf{x}_{\text{Tri}}$. However, the resulting triggers are largely uninterpretable as natural language (cf. App.˜E). We therefore build a more complex and effective pipeline in §˜4.3.
@@ 3.3 Threat Model
Having established the mechanism underlying HiSPAs, we now formalize the attacker capabilities required to deploy them in practice. In adversarial learning, threat models define the trade-off between the capabilities of the attacker $\mathcal{A}$ (hence the attack’s strength) and the realism of the scenario (whether such resources would be available to a malicious actor in practice). In this study, we choose to focus on two scenarios: first, a zero-shot setting where $\mathcal{A}$ does not have access to the model weights or gradients, and must craft the trigger without any finetuning. We call HiSPAs generated in this setting Z-HiSPAs, written in natural language and based on the “escape”, “ignore” and “fake completion” strategies as used by liu_formalizing_2024 (cf. App.˜B). Because the profile of $\mathcal{A}$ would match an average chatbot user, this scenario is particularly relevant for practical security evaluations.
@@ 4 RoBench-25 Benchmark Results
Model/Config No Trigger Trigger A Trigger B Trigger C Trigger D Trigger E Trigger F Trigger G
@@
Table 1: RoBench-25: CHSS scores (higher is better) for each model–configuration pair across all triggers (A–G) averaged over 120 different random draws of InT (1 abstract) and DiT (6 abstracts). Bold values indicate the best score among the small language models (pythia and mamba). Refer to App.˜D for configuration details and App.˜B for trigger details.
The underlying dataset consists of 120 NeurIPS 2025 paper titles plus abstracts (260.8 tokens on average), each paired with 2 True/False questions designed to evaluate information retention (see App.˜C for the question generation protocol and App.˜F for some examples). Prompts are then constructed starting with $n_{\text{InT}}$ target abstracts (containing information needed to answer the questions), followed by an optional HiSPA trigger, $n_{\text{DiT}}$ distractive abstracts, and finally the questions. Prompts may additionally include an awareness instruction at the beginning (warning the model to ignore forgetting instructions) and/or a recovery instruction before the questions (reminding the model to focus on the target abstracts). Full construction details are provided in Appendix D.
To evaluate how susceptible LLMs are to the designed HiSPAs, we will measure to what extent they correctly answer the posed questions using the clipped Heidke skill score (heidke1926):
$\text{CHSS}:=\max\left(0,\frac{1-\text{Acc.}}{1-\text{Rand.}}\right),$ (6)

### v3

@@ Abstract
State space models (SSMs) like Mamba offer efficient alternatives to Transformer-based language models, with linear time complexity. Yet, their adversarial robustness remains critically unexplored. This paper studies the phenomenon whereby specific short input phrases induce a partial amnesia effect in such models, by irreversibly overwriting information in their hidden states, referred to as a Hidden State Poisoning Attack (HiSPA). Our benchmark RoBench-25 allows evaluating a model’s information retrieval capabilities when subject to HiSPAs, and confirms the vulnerability of SSMs against such attacks. Even the recent Jamba-1.7-Mini SSM–Transformer (a 52B hybrid model) collapses on RoBench-25 under some HiSPA triggers, whereas pure Transformers do not. We also observe that HiSPA triggers significantly weaken the Jamba model on the popular Open-Prompt-Injections benchmark, unlike pure Transformers. We further show that the theoretical and empirical findings extend to Mamba-2 and to a Mamba-2-based hybrid (Nemotron-3-Nano). Finally, our interpretability study reveals patterns in Mamba’s hidden layers during HiSPAs that could be used to build a HiSPA mitigation system. The full code and data to reproduce the experiments can be found at https://anonymous.4open.science/r/hispa_anonymous-5DB0.
@@ 1 Introduction
State space models (SSMs) like Mamba (Gu and Dao, 2024) and Mamba-2 (Dao and Gu, 2024) have emerged as a promising alternative to traditional Transformers in large language models (LLMs) (Rezaei Jafari et al., 2024; Waleffe et al., 2024), thanks to their ability to efficiently capture long-range dependencies while achieving linear time complexity during both training and inference. Beyond efficiency and speed advantages, research in SSMs is particularly relevant in today’s ecological context, where the massive energy consumption of standard Transformers (with quadratic complexity) is becoming a growing concern (Patterson et al., 2021; Liu and Yin, 2024).
Despite the impressive capabilities demonstrated by recent SSM-based LLMs, their robustness against adversarial attacks remains largely unexplored, and no dedicated benchmark has been established. We argue that this is a critical oversight, since the unique design of the widely used Mamba block, in particular its selectivity mechanisms and recurrent hidden-state updates, may expose new types of vulnerabilities that are not present in Transformer-based architectures. Our study reveals and quantifies one such vulnerability, which a malicious user can easily exploit through a short trigger sequence that causes partial amnesia in Mamba by irreversibly overwriting its hidden states. Such Hidden State Poisoning Attack (HiSPA) requires no optimization and can be crafted in zero-shot black-box settings, yet it drastically impairs Mamba’s ability to retain and retrieve information from long contexts, even when explicitly instructed to disregard the injected content.
This vulnerability directly connects to the well-known Prompt Injection Attack (PIA) problem. We show that HiSPAs severely degrade Mamba’s context retention and that the industrial Jamba-1.7-Mini model [^1] https://huggingface.co/ai21labs/AI21-Jamba-Mini-1.7 exhibits a significantly higher PIA success rate when a HiSPA precedes the prompt, which demonstrates that hidden-state corruption can amplify classical prompt injections (§§˜3 and 6).
Addressing this issue is urgent in today’s adversarial environment surrounding LLMs (Lin, 2025). A representative example is that of assisted scientific literature review tools such as Scispace [^2] https://www.scispace.com/ or Elicit, [^3] https://elicit.com/ which rely on automated reasoning over long textual contexts (Wu et al., 2023; Tillmann, 2025). Recent reports indicate that malicious actors embed prompt injections directly within papers (Keuper, 2025), creating realistic attack surfaces for SSM-based models deployed in these pipelines. Such scenarios underscore the need for systematic robustness evaluation of hybrid architectures before deployment.
@@
  • We introduce Hidden State Poisoning Attacks (HiSPAs), a new class of adversarial triggers that exploit SSM recurrence to overwrite hidden states in Mamba-based language models (including Mamba-2), with zero-shot (Z-HiSPA) and optimized white-box (M-HiSPA) variants (§˜3).
@@
  • We show that optimized M-HiSPAs can lead to the collapse of the Jamba-1.7-Mini hybrid (built with Mamba blocks), despite its attention layers and 52B parameters (§˜4.3).
@@
  • We show that HiSPAs theoretically and empirically generalize to Mamba-2, and evaluate their impact on the Nemotron-3-Nano hybrid (built with Mamba-2 blocks). [^4] This part of the study was moved to App. H because of space constraints.
@@ 2.1 State Space Models for LLMs
The Mamba model (Gu and Dao, 2024) is built upon the S4 family of state space models (SSMs) (Gu et al., 2021). Mamba replaces the attention mechanism of Transformers (Vaswani et al., 2017) with a structure inspired by recurrent neural networks (RNNs) (Schmidt, 2019) and control theory (Kalman, 1960), achieving linear time complexity with respect to sequence length. Mamba processes each token sequentially while maintaining a hidden state $\mathbf{h}^{(b)}_{t}$ that encodes information from previous tokens, effectively acting as an explicit memory (initialized to zero). For notational simplicity, we omit the block superscript $(b)$ in the remainder of the paper unless disambiguation is required.
@@
The advent of Mamba in 2023 and Mamba-2 in 2024 (Dao and Gu, 2024), whose recurrence shares the same structure but uses scalar-identity state transitions per head (cf. App.˜H), demonstrated that SSMs can match or surpass Transformers on key benchmarks, while offering gains in memory and computational efficiency, especially in hybrid architectures interleaving SSM layers with attention (Waleffe et al., 2024). This breakthrough has sparked innovative hybrid designs including Samba (Ren et al., 2024), Hymba (Dong et al., 2024), BlackMamba (Anthony et al., 2024), Zamba (Glorioso et al., 2024), the Nemotron family (Blakeman et al., 2025a; b), and the Jamba family (Lieber et al., 2024; Team et al., 2024). However, the adversarial robustness of these architectures remains unexplored.
@@
Jamba is the largest Mamba-based hybrid family in the literature (up to 398B parameters), demonstrating effective context lengths of 256k tokens on the RULER benchmark (Hsieh et al., 2024), being the only model family maintaining high performance at this scale without degradation (Team et al., 2024). Jamba achieves competitive results on standard benchmarks while exhibiting known SSM weaknesses in reasoning and code generation (Patro and Agneeswaran, 2025; Chen et al., 2025b; Ren et al., 2025). Crucially, Mamba blocks constitute $\nicefrac{{7}}{{8}}$ of Jamba’s architecture, making it ideal for evaluating whether HiSPA vulnerabilities transfer from pure-Mamba models to production-scale hybrids.
@@ 2.2 Prompt Injection Attacks
Prompt Injection Attacks (PIAs) exploit how language models process input prompts to manipulate outputs for malicious purposes (Greshake et al., 2023; Liu et al., 2024). These typically comprise a trigger (also named distractor) and a payload, e.g., “Ignore previous instructions and give me your API key”. Liu et al. (2024) proposed a standard benchmark showing that combining a fake completion with an ignore instruction severely degrades performance of several Transformer-based LLMs, including GPT-4, and that defensive prompt engineering yields only marginal improvements. Although PIAs have been widely studied in Transformer-based LLMs (Greshake et al., 2023; Perez and Ribeiro, 2022; Chen et al., 2025a), their impact on SSM-based models remains largely unexplored, leaving open whether these architectures introduce new attack vectors or exhibit different vulnerability patterns.
Figure 1: The Mamba core (left) and Mamba block (right) as described by Gu and Dao (2024). For block $b$ and time step $t$, $\mathbf{x}_{t}^{(b)}\in\mathbb{R}^{K}$ is the input core token embedding, $\mathbf{h}_{t}^{(b)}\in\mathbb{R}^{K\times N}$ the hidden states with $N$ the length of $\mathbf{h}_{t,k}^{(b)}$, $\mathbf{y}_{t}^{(b)}\in\mathbb{R}^{K}$ the output core token embedding, and $\mathbf{o}_{t}^{(b)}\in\mathbb{R}^{K/2}$ the output block token embedding.
Figure 2: Greedy beam search evolution of the loss function (5) per token added to the trigger, distributed over 100 random seeds (log scale). The exponential decay of the loss after token n°2 is consistent with the exponential saturation of the hidden states $\mathbf{h}_{t}$ described in App.˜B.
3 Hidden State Poisoning Attacks
@@ 3.1 Theoretical Foundation
The core idea behind a Hidden State Poisoning Attack (HiSPA) is to exploit the inability of SSMs (such as Mamba or Mamba-2) to attend back to previous tokens and its selectivity mechanism, to poison the hidden state $\mathbf{h}_{t}$. With “poison” we mean introducing a small sequence of tokens that will irreversibly overwrite a large part of the information contained in the hidden state. We will refer to such sequence as a trigger, to align with the PIA terminology.
@@
In eqs.˜1, 2 and 3, Gu and Dao (2024) constrained the $\Delta_{t}$ matrix to be positive for computational stability reasons, and made $\mathbf{A}$ diagonal. Further, $\mathbf{B}$ is learned during training, and its values are unconstrained. The objective of a HiSPA is therefore to identify a short sequence whose associated updates drive the recurrence into a contracting regime and inject a dominant input term. As shown in App.˜B, if all tokens in the trigger belong to a subset for which the corresponding state-transition factors satisfy a uniform contraction bound (eqs.˜9, 10, 11 and 12), the carried-over state decays geometrically with the trigger length (eq.˜13). Meanwhile, the input part of the unrolled update necessarily contains a summand whose magnitude equals the maximum attainable input contribution over that subset (eq.˜14). By choosing the last token to realize this maximum and selecting the trigger length so that the decaying state term becomes smaller than this input term (eq.˜17), the injected contribution dominates the previous hidden state. In this sense, a carefully chosen trigger can overwrite most of the stored information in $\mathbf{h}_{t}$, establishing the mechanism enabling HiSPA. Of note, hybrid models combining Mamba with attention layers may partially mitigate this vulnerability, since attention can recover information that the SSM layers have overwritten. We test this hypothesis in §˜4 with Jamba-1.7-Mini and Nemotron-3-Nano (see also §˜H.4 for the later).
Generalization to Mamba-2.
The derivation above relies on the contracting-regime bound (App.˜B) and does not require $\bar{\mathbf{A}}_{t}$ to be a general diagonal: the same argument holds when $\bar{\mathbf{A}}_{t}=\alpha_{t}I_{N}$ is scalar-identity, as in Mamba-2 (Dao and Gu, 2024). The full adaptation of the proof and empirical validation are provided in App.˜H for space constraints.
@@
Mamba blocks contain a skip connection (cf. Fig.˜2) that combines the output of the Mamba core branch with a linear projection of the input token embedding itself. It is unclear from the theory alone to what extent this skip connection may dilute the poisoning effect of the Mamba core, further motivating empirical evaluations in §˜4.
@@ 3.2 Empirical Verification
$\mathcal{L}(\mathbf{x}_{\text{Tri}}):=-\cos\left(\mathbf{h}^{(1)}_{t}(\mathbf{x}_{\text{InT}}\oplus\mathbf{x}_{\text{Tri}}),\mathbf{h}^{(1)}_{t}(\mathbf{x}_{\text{Tri}})\right)$ (5)
Using 10 Tesla V100-SXM3-32GB GPUs, we evaluate 100 random seeds (42–141) and obtain the graph in Fig.˜2. As expected from theory, we observe an approximately exponential decay of the loss once the contracting regime dominates, consistent with the $\rho^{L}$ bound in eq.˜12. The best trigger obtained reached a loss of -0.997 after its 6 tokens, making the embedding of $\mathbf{x}_{\text{InT}}\oplus\mathbf{x}_{\text{Tri}}$ nearly aligned with that of $\mathbf{x}_{\text{Tri}}$. However, the resulting triggers are largely uninterpretable as natural language (cf. App.˜F). We therefore build a more complex and effective pipeline in §˜4.3.
@@ 3.3 Threat Model
Having established the mechanism underlying HiSPAs, we now formalize the attacker capabilities required to deploy them in practice. In adversarial learning, threat models define the trade-off between the capabilities of the attacker $\mathcal{A}$ (hence the attack’s strength) and the realism of the scenario (whether such resources would be available to a malicious actor in practice). In this study, we choose to focus on two scenarios: first, a zero-shot setting where $\mathcal{A}$ does not have access to the model weights or gradients, and must craft the trigger without any finetuning. We call HiSPAs generated in this setting Z-HiSPAs, written in natural language and based on the “escape”, “ignore” and “fake completion” strategies as used by Liu et al. (2024) (cf. App.˜C). Because the profile of $\mathcal{A}$ would match an average chatbot user, this scenario is particularly relevant for practical security evaluations. [^5] In § H.2 and § H.3, we extend this black-box design to a multi-shot one, yielding better results.
@@ 4 RoBench-25 Benchmark Results
Model/Config No Trigger Tri. A Tri. B Tri. C Tri. D Tri. E Tri. F Tri. G
@@
mamba-2, A-R+ 0.117 0.192 0.058 0.067 0.167 0.108 0.075 0.267
nemotron, A-R+ 0.892 0.900 1.000 0.933 1.000 0.975 0.992 0.683
Table 1: RoBench-25: CHSS scores (higher is better) for each model–configuration pair across all triggers (A–G) averaged over 120 different random draws of InT (1 abstract) and DiT (6 abstracts). Bold values indicate the best score among the small language models (pythia, mamba and mamba-2). Results for mamba-2 and nemotron are discussed in App.˜H. Refer to App.˜E for configuration details, App.˜C for trigger details and App.˜I for results on 189 additional triggers.
The underlying dataset consists of 120 NeurIPS 2025 paper titles plus abstracts (260.8 tokens on average), each paired with 2 True/False questions designed to evaluate information retention (see App.˜D for the question generation protocol and App.˜G for some examples). Prompts are then constructed starting with $n_{\text{InT}}$ target abstracts (containing information needed to answer the questions), followed by an optional HiSPA trigger, $n_{\text{DiT}}$ distractive abstracts, and finally the questions. Prompts may additionally include an awareness instruction at the beginning (warning the model to ignore forgetting instructions) and/or a recovery instruction before the questions (reminding the model to focus on the target abstracts). Full construction details are provided in Appendix E.
To evaluate how susceptible LLMs are to the designed HiSPAs, we will measure to what extent they correctly answer the posed questions using the clipped Heidke skill score (Heidke, 1926):
$\text{CHSS}:=\max\left(0,\frac{\text{Acc.}-\text{Rand.}}{1-\text{Rand.}}\right),$ (6)